OpenStudy (anonymous):

Hi! Good Afternoon. Can someone help me about the area of polar coordinates? I saw in a certain example that the interval Of the equation r=2-sinΘ and r=3sinΘ is Π/6≤Ɵ≤Π/2. Can someone explain how it becomes?

OpenStudy (aravindg):

hey welcome to openstudy !!

OpenStudy (anonymous):

Can you answer my question. Thank you

zepdrix (zepdrix):

The inverval of the equation? Hmm I'm not quite sure what that means. Here is a nice pretty graph of the two functions. https://www.desmos.com/calculator/xlhjges48e It seems they intersect at \(\large \theta=\pi/6\) and again at \(\large \theta=5\pi/6\).

OpenStudy (anonymous):

yes. it intersects at θ=π/6 and again at θ=5π/6, but the interval that i posted is also right? is it? then i will just multiply it by 1/2? am i right?

zepdrix (zepdrix):

What are you trying to do exactly? Find the area of the intersecting regions? If so, then yes you could do that. Setup your interval from \(\large \pi/6\) to \(\large \pi/2\) and just multiply the by \(\large 2\) since the area is symmetric up there.

OpenStudy (anonymous):

yes, that was exactly what i mean.. but, is there any formula to find the intersecting regions?

zepdrix (zepdrix):

For finding points of intersection? \[\large r=2-\sin \theta \qquad \qquad r=3\sin \theta\] Set your two equations equal to one another,\[\large 2-\sin \theta=3\sin \theta\] And then solve for \(\large \theta\). We will do so by first solving for \(\large \sin\theta\) and from there remembering our special angles. Add sin theta to both sides,\[\large 2=4\sin \theta\]Divide by 4,\[\large \frac{1}{2}=\sin\theta\] Can you recall what special angle produces 1/2 from the sine function? :D

zepdrix (zepdrix):

Those are special angles on the unit circle c: \(\large \sin \dfrac{\pi}{6}=\dfrac{1}{2}\) \(\large \sin \dfrac{5\pi}{6}=\dfrac{1}{2}\) You should definitely have the unit circle memorized if you're doing this for a school class c: Yah you might be able to get the angle by arcsine, if you have a good calculator it might even give it to you as a nice fraction instead of a messy decimal.

OpenStudy (anonymous):

okay i think i remember those special angles. but i did not memorize it.. by the way, can you give me an example that i can practice to solve at home afterwards? if its okay. :D